# Shaping the Safety Boundaries: Understanding and Defending Against Jailbreaks in Large Language Models

### v1

@@ Abstract
Jailbreaking in Large Language Models (LLMs) is a major security concern as it can deceive LLMs to generate harmful text. Yet, there is still insufficient understanding of how jailbreaking works, which makes it hard to develop effective defense strategies. We aim to shed more light into this issue: we conduct a detailed large-scale analysis of seven different jailbreak methods and find that these disagreements stem from insufficient observation samples. In particular, we introduce safety boundary, and we find that jailbreaks shift harmful activations outside that safety boundary, where LLMs are less sensitive to harmful information. We also find that the low and the middle layers are critical in such shifts, while deeper layers have less impact. Leveraging on these insights, we propose a novel defense called Activation Boundary Defense (ABD), which adaptively constrains the activations within the safety boundary. We further use Bayesian optimization to selectively apply the defense method to the low and the middle layers. Our experiments on several benchmarks show that ABD achieves an average DSR of over 98% against various forms of jailbreak attacks, with less than 2% impact on the model’s general capabilities.
@@
Lang Gao1,2,  Xiangliang Zhang3, Preslav Nakov1, Xiuying Chen1222Corresponding author. 1MBZUAI  2Huazhong University of Science and Technology 3University of Notre Dame {Lang.Gao, Preslav.Nakov, Xiuying.Chen}@mbzuai.ac.ae xzhang33@nd.edu
@@ 1 Introduction
The widespread use of Large Language Models (LLMs) across various fields Kaddour et al. (2023) has raised concerns about the safety of the output and the robustness of the model. It has been shown that jailbreak attacks, which use crafted prompts to deceive LLMs into generating harmful content, can bypass LLM’s safety alignment Liu et al. (2024b); Zou et al. (2023), and a lot of research has focused on developing defense mechanisms and counter-prompts to mitigate such attacks Robey et al. (2023); Xie et al. (2024).
@@
Understanding the internal mechanism of how jailbreaks work is crucial for improving defenses and enhancing safety measures. However, studies explaining jailbreaks remain largely ambiguous and controversial Yu et al. (2024); Ball et al. (2024a); Lin et al. (2024). From an explainability perspective, debates revolve around which layers are most responsible for a successful attack: He et al. (2024) prioritize the low layers, Zhou et al. (2024); Shen et al. (2024) focus on the middle layers, while Li et al. (2024) highlight the deep layers. Another contradiction lies in whether activations gradually shift as layers deepen or abruptly transition in the middle layers Zhao et al. (2024a); Shen et al. (2024). Our experiments suggest that these inconsistencies may stem from the limited sample size of prompts, typically around 100, used in prior analyses. From a defense perspective, existing approaches often rely on additional training processes Zhao et al. (2024a) or probing a limited number of samples Shen et al. (2024), which affects the utility of LLMs.
@@
Here we aim to better understand the mechanism of how jailbreaking works. In particular, we provide our explanation of jailbreaks based on a comprehensive analysis of over 30,000 samples, a significantly larger scale than previous studies. As an example, Figure 1 shows the two-dimensional projection of three types of prompts: benign prompts that contain no harmful information, harmful prompts that attempt to induce the LLM to generate harmful content but fail, and various jailbreak prompts that successfully induce harmful outputs, collected from different types of jailbreak attack methods. The projection is based on the last token representations across different layers, referred to as activations, as they capture the model’s overall understanding of the entire input sequence Radford et al. (2019); Zou et al. (2023). We can see that different prompts exhibit clustering effects, where the jailbreak activations are distinct from the activation space of harmful prompts. This indicates that in each layer, the LLM performs a rigorous safety check within a safety boundary, effectively containing most of the harmful activations within this controlled space. Jailbreak prompts, however, bypass these safety checks by shifting activations beyond the safety boundary into an alternative, unmonitored activation space. Notably, the extent of this shift varies across different layers of the LLM, with the low and the middle layers exhibiting the most pronounced effects. This suggests that these layers are critical to the success of jailbreaks.
@@
This finding motivates a new defense mechanism that we propose, Activation Boundary Defense (ABD), which constrains jailbreak prompt activations within a safety boundary by applying penalties to activation values. Specifically, we impose minimal penalties within the boundary but sharply increase them beyond it, thereby preserving model utility. The constraint function considers various factors, including the defense layer and the defense dimension, which balance performances on defense and general tasks. We propose a novel adaptive objective that maximizes the Defense Success Rate (DSR) while minimizing the number of penalized layers. This objective is implemented with an adapted Bayesian Optimization (BO) method Jones et al. (1998); Mockus (1974), which iteratively proposes and refines selection suggestions.
@@
Experiments on AdvBench Zou et al. (2023) show that ABD has good generalization, achieving an average DSR of over 98% against various forms of jailbreak attacks. In the general ability tasks, the performance drop of the model equipped with ABD does not exceed 2%, which is significantly lower compared to other defenses, such as 37% for Retokenization Jain et al. (2023). This demonstrates that our method has a minimal impact on the model’s overall abilities. In summary, our contributions are threefold:
@@
  • We uncover a comprehensive activation distribution and introduce the safety boundary, resolving contradictions in prior work and highlighting the vulnerability of the low and the middle layers.
@@
  • We propose a lightweight, extensible defense that penalizes only targeted samples and a few key layers, ensuring efficiency and precision.
@@
  • Our experiments on benchmark datasets demonstrate that ABD achieves an average DSR of over 98% against various jailbreak attacks, with less than a 2% impact on the general capabilities.
@@ 2 Related Work
Jailbreak attacks craft prompts to induce harmful outputs from LLMs. Early approaches, like DAN Shen et al. (2023), rely on manually designed prompts. Later research proposed automatic methods. Optimization-based methods Zou et al. (2023); Liu et al. (2024b); Guo et al. (2024) design optimization strategies to iteratively edit and refine original harmful prompts, enhancing their stealthiness. Model-based methods Chao et al. (2023); Ding et al. (2024); Paulus et al. (2024) utilize an attacker LLM to autonomously query and refine jailbreak prompts. For rule-based jailbreak methods, Jiang et al. (2024) and Liu et al. (2024c) rewrite prompts using fixed rules that are sufficiently aggressive to deceive the model. In this work, we aim to analyze the mechanisms of these jailbreak methods to develop more effective defense strategies.
@@
Unlike jailbreak methods, defense strategies enhance model safety and robustness by reformulating inputs to counter jailbreak prompts, such as backtranslation Wang et al. (2024), paraphrasing Jain et al. (2023), and reminding Xie et al. (2023). Some researchers use indicators to distinguish jailbreak prompts, e.g., classify input sequences based on perplexity and sequence length Alon and Kamfonas (2023), or classify jailbreak activations using linear classifiers to control whether to answer the query Zhao et al. (2024a). A new trend in defense involves directly manipulating the representations of the model Shen et al. (2024); Xu et al. (2024); Li et al. (2024); Liu et al. (2024a) and editing the model Zhao et al. (2024b), which is more efficient and cost-effective. Unlike these methods, ours avoids extra tokens or modules by directly constraining activations within a safety boundary, ensuring greater efficiency and simplicity.
@@
The growing concern about LLM safety has sparked increasing interest in interpreting LLM features in jailbreak prompts. For example, Ball et al. (2024b) identified a common mechanism whereby jailbreaks reduce the harmfulness perception in most LLMs. Similarly, Li et al. (2024) investigated patterns that trigger the model to recognize safety issues. Zhou et al. (2024) discovered that the vocabulary mappings of activations significantly changed when processing jailbreak inputs. Research efforts have also focused on proposing corresponding defense methods based on interpretability results. For instance, Zhao et al. (2024a) and Shen et al. (2024) model how jailbreak activations transfer between benign and harmful activation spaces as layers deepen. They designed adaptive defenses whose strength varies across different layers. However, the limited data in previous studies often led to controversy and ambiguity. In contrast, we leverage over 30,000 samples to better understand jailbreaks and design effective defenses.
@@ 3.1 Controversy in Literature
What are the key layers?
There have been different opinions about which layers are most important, mainly due to varying perspectives. For example, He et al. (2024) argued that low layers are essential because they treat jailbreak activations as benign ones. Zhou et al. (2024) believe that low and middle layers are essential based on tokens generated from activations. Li et al. (2024) argued that deep layers are critical, as the concept they proposed, called safety pattern, has greater values in these layers. Among these works, the first work similarly focuses on activations, as does our paper.
Upon examining their experimental setup, we identified a major limitation: the small number of samples used in the analysis. As shown in Figure 2(a), before scaling the data, benign and harmful activations are linearly separable, with most jailbreak samples being misclassified as benign, which is consistent with the findings from He et al. (2024). However, after scaling the data, half of the jailbreak activations shift to align with harmful ones, and the separation becomes nonlinear. This highlights the risk of misleading from small-scale studies and underscores the importance of large-scale analysis.
@@
There are also different perspectives on the internal mechanism of jailbreak. Specifically, Zhao et al. (2024a) viewed jailbreak as a gradual process where activations transition from harmful to benign spaces from low to deep layers, while Shen et al. (2024) argued that jailbreak abruptly manifests in deeper layers, with activations positioned between harmful and benign spaces.
Figure 2: PCA visualization of the activations highlights how conclusions are drawn from tiny sample sizes, such as the side on which the jailbreak activation lies, may not hold with larger datasets. (a) Top: 100 samples per type (following He et al. (2024)); Bottom: Benign and harmful samples increased to 5000. (b) Top: 60 samples per type (similar to Zhao et al. (2024a)); Bottom: All samples increased to 500.
This work used no more than 100 samples per activation type, and we identify a similar issue to the one discussed earlier. Figure 2(b) shows that before scaling the data, the jailbreak activations are aligned to benign activations, which is consistent with findings from Zhao et al. (2024a); while after scaling the data, jailbreaking activations no longer align with either harmful or benign activations. When compared to Shen et al. (2024), similar conflicting perspectives are observed when scaling up the data. Further details can be found in Appendix C.1. These conflicting perspectives form the basis of the jailbreak mechanisms in these works; their conclusions — such as whether activations gradually shift or abruptly transition — are also questionable.
@@ 3.2 Experimental Settings
Dataset.
Our dataset consists of 32,507 samples in three categories: benign, harmful, and jailbreak. For benign samples, we used a subset from the Alpaca dataset Taori et al. (2023), as it has been carefully curated to ensure that only safe content is included. For harmful samples, we collected five datasets containing harmful queries from RedEval Bhardwaj and Poria (2023) and AdvBench Zou et al. (2023). These datasets encompass a wide variety of harmful prompts designed to elicit unsafe outputs. For jailbreak samples, we applied seven different jailbreak attack methods, including GCG-Individual and GCG-Universal Zou et al. (2023), AdvPrompter Paulus et al. (2024), COLD-Suffix Attack Guo et al. (2024), AutoDAN Liu et al. (2024b), PAIR Chao et al. (2023), and ArtPrompt Jiang et al. (2024), on AdvBench. Detailed statistics can be found in Appendix A.
Model.
We used Vicuna-7B-v1.3 Chiang et al. (2023), a 32-layer Transformer model, as the LLM to study. It is tuned upon a safety-aligned LLM and is able to identify certain harmful prompts. However, as demonstrated in previous work Chu et al. (2024), it can also be easily deceived and generate harmful information during jailbreak attacks. This makes it an excellent example for studying both attack mechanisms and defense strategies.
@@
Our projection was based on the activation, the last token vector of the input, as it captures the model’s overall understanding of the entire input sequence Radford et al. (2019); Zou et al. (2023). As discussed in §3.1, linear classification may be insufficient to model the boundary between different prompts. Therefore, we adopted Multi-Dimensional Scaling (MDS) Carroll and Arabie (1998) as our dimensionality reduction method, as MDS is more effective than PCA in handling large-scale and complexly distributed data Anowar et al. (2021). Moreover, it better captures the global distribution structure compared to t-SNE Van der Maaten and Hinton (2008).
@@ 3.3.1 Activation Distribution
We present the activation distributions of representative layers in Figure 1, with the full version available in Appendix B. Our key observations are as follows: (1) Harmful and benign activations overlap and are not linearly separable in most layers, especially in low and middle layers, as shown in Figure 1(a) and Figure 1(b), where benign and harmful activations are overlapping with no clear linear decision boundaries to separate them. This highlights the intricate interplay between harmful and benign activations, indicating that forcing jailbreak activations into a linear benign-vs-harmful classification is impractical. (2) Jailbreak activations significantly shift out of the harmful activation space in most layers, forming an independent region with low overlap with harmful activations. This pattern is consistent across various jailbreaks, suggesting that jailbreaks achieve deception by shifting activations. Such shift is evident in all layers, indicating that jailbreaks begin in low layers and continue to affect the model across all layers, rather than starting in middle or deep layers Zhou et al. (2024); Shen et al. (2024). Moreover, the shift is the most pronounced in the low and middle layers, implying their importance in jailbreaks.
@@ 3.3.2 Safety Boundary
Based on the observation of jailbreak and harmful activations, we propose the following hypothesis: A safety boundary exists that ensures sensitivity to harmful content within its activation space, but jailbreak prompts can bypass it, generating harmful responses. To verify the existence of the safety boundary, we conduct experiments and present two concepts described below.
We first explore Randomized Activation Shifting (RAS), an approximation of jailbreak attacks that introduces a random value shift to the original activation. Specifically, assuming the activation on the $l$-th layer is $a_{l}$, we shift it by a distance $r$ in a random direction $\hat{u}$:
@@
Figure 3(a) shows the changes in Defense Success Rate (DSR) under different values of $r$ for four selected layers. The definition of this metric is provided in the Appendix. Initially, when $r=0$, the corresponding DSR is over 99%, as the harmful text has activations within the safety boundary. Then, as the distance $r$ increases, the DSR decreases for all layers, indicating that harmful activations become harder for the model to detect. Notably, regardless of the selected layer, a steep drop in DSR is observed at a specific range (e.g., $r=10-20$ on layer 0), while the decline is more moderate in other regions. This observation suggests the presence of a safety boundary.
For more accurate boundary detection, we provide the second visualization, which introduces a measurable concept called Most Vulnerable Distance (MVD) as an approximation of the boundary. MVD represents the specific distance at which the DSR experiences the steepest drop, indicating the point where the model’s ability to detect harmful activations becomes most vulnerable:
@@
where $\frac{\mathrm{d(DSR)}}{\mathrm{d}r}$ denotes the rate of change of DSR with respect to $r$. As shown in Figure 3(b), MVD increases with the depth of the model’s layers. This implies that the safety boundary also expands in deeper layers, requiring more substantial shifts to compromise the model’s safety checks. The findings align closely with our intuitive observations in Figure 1. Figure 1(a) shows that in low layers, jailbreak activations are shifted in various directions, but all evidently surpass the safety boundary. In the middle layers presented in Figure 1(b), most jailbreak activations are beyond the safety boundary, and they are more aligned with benign activations. In the deep layers shown in Figure 1(c), jailbreak activations cluster within a region, which overlaps a little with the safety boundary.

### v2

@@ Abstract
Jailbreaking in large language models (LLMs) poses major security risks by tricking models into generating harmful text. However, there is limited understanding of how jailbreaking operates, which makes it difficult to develop effective defenses. In this work, we conduct a large-scale analysis of seven jailbreak methods and uncover that inconsistencies in previous studies arise from insufficient observation samples. Our analysis reveals that jailbreaks shift harmful activations beyond a defined safety boundary, where LLMs become less sensitive to harmful information. We also find that the low and the middle layers are critical in driving these shifts, while deeper layers play a lesser role. Leveraging these insights, we propose a novel defense mechanism called Activation Boundary Defense (ABD), which adaptively constrains activations within the safety boundary. To further optimize performance, we use Bayesian optimization to select the most effective layers for ABD application and confirm that the low and middle layers have the greatest impact, consistent with our earlier observations. Experiments across multiple benchmarks demonstrate that ABD achieves an average defense success rate (DSR) of over 98% against various jailbreak attacks, with less than 2% impact on the model’s overall capabilities.
@@
Lang Gao1,2,  Jiahui Geng1,  Xiangliang Zhang3, Preslav Nakov1, Xiuying Chen1thanks: Corresponding author. 1MBZUAI  2Huazhong University of Science and Technology 3University of Notre Dame {Lang.Gao, Jiahui.Geng, Preslav.Nakov, Xiuying.Chen}@mbzuai.ac.ae xzhang33@nd.edu
[…]
WARNING: This paper contains model outputs that may be considered offensive.
@@ 1 Introduction
The widespread use of Large Language Models (LLMs) across various fields Kaddour et al. (2023); Chen et al. (2025); Xie et al. (2025); Liu et al. (2024c) has raised concerns about the safety of the output and the robustness of these models Ye et al. (2024); Huang et al. (2025); Wang et al. (2025). It has been shown that jailbreak attacks, which use crafted prompts to deceive LLMs into generating harmful content, can bypass LLM’s safety alignment Liu et al. (2024b); Zou et al. (2023); Song et al. (2025), and a lot of research has focused on developing defense mechanisms and counter-prompts to mitigate such attacks Robey et al. (2023); Xie et al. (2024).
@@
Understanding the internal mechanisms of jailbreak attacks is crucial for both explaining their operation and developing effective defenses. While existing studies have made initial explorations Yu et al. (2024); Ball et al. (2024a); Lin et al. (2024), fundamental disagreements remain about the attack mechanisms. The academic community continues to debate two key aspects of explainability: (i) The localization of attack impacts across the model layers, with competing findings suggesting primary responsibility in low He et al. (2024), middle Zhou et al. (2024b); Shen et al. (2024), or deep layers Li et al. (2024), and (ii) the progression pattern of malicious activation, with conflicting observations about gradual shifts versus abrupt transitions in the middle layers Zhao et al. (2024a); Shen et al. (2024). Our experiments suggest that these inconsistencies may stem from methodological limitations in prior analysis, particularly their reliance on small prompt samples (typically $\sim 100$). On the defense front, current approaches face dual challenges: Most detection methods either require additional training Zhao et al. (2024a) or demonstrate limited generalizability due to small-sample probing Shen et al. (2024), while mitigation strategies often degrade normal model capabilities.
@@
In this work, we aim to better understand the mechanism of how jailbreaking works. In particular, we provide our explanation of jailbreaks based on a comprehensive analysis of over 30,000 samples, a significantly larger scale than previous studies. Figure 1 shows the projection of three types of prompts: benign prompts that contain no harmful information, harmful prompts that attempt to induce the LLM to generate harmful content but fail, and various jailbreak prompts that successfully induce harmful outputs, collected from different types of jailbreak attack methods. The projection is based on the last token representations across different layers, referred to as activations, as they capture the model’s overall understanding of the entire input sequence Radford et al. (2019); Zou et al. (2023). Our results reveal distinct clustering patterns, where different prompts form separate activation clusters. Notably, harmful prompts form a contained cluster within a specific region that we define as the safety boundary, representing the LLM’s internal safety check that typically prevents harmful activations from escalating into harmful outputs. In contrast, jailbreak prompts push activations beyond the safety boundary into an unmonitored space where harmful content can be generated. This shift is most pronounced in the low and middle layers, indicating their critical role for jailbreak success.
@@
This finding motivates a new defense mechanism that we propose, Activation Boundary Defense (ABD), which constrains jailbreak prompt activations within a safety boundary by applying penalties to activation values. Specifically, we impose minimal penalties within the boundary, but sharply increase them beyond it, thereby preserving model utility. The penalty function considers various factors, including the defense layer and the defense dimension, which balance the performance on defense and on general tasks. We propose a novel objective that maximizes the Defense Success Rate (DSR) while minimizing the number of penalized layers. This objective is implemented with an adapted Bayesian Optimization method Jones et al. (1998), which iteratively proposes and refines selection suggestions.
@@
Experiments on AdvBench Zou et al. (2023) show that ABD has good generalization, achieving an average DSR of over 98% against various forms of jailbreak attacks. For general ability tasks, the performance drop of the model equipped with ABD does not exceed 2%, which is substantially lower than the drop of up to 37% with other defenses.
Our contributions can be summarized as follows:
@@
$\bullet$ We uncover a comprehensive activation distribution and introduce the safety boundary, resolving contradictions in prior work and highlighting the vulnerability of the low and middle layers.
@@
$\bullet$ We propose a lightweight, extensible defense that penalizes only targeted samples and a few key layers, ensuring efficiency and precision.
@@
$\bullet$ Our experiments on benchmark datasets demonstrate that ABD achieves an average DSR of over 98% against various jailbreak attacks, with less than a 2% impact on general capabilities.
@@ 2 Related Work
Jailbreak attacks craft prompts to induce harmful outputs from LLMs, evolving from early manually designed approaches to more sophisticated techniques. Early methods, such as DAN Shen et al. (2023), rely on manually constructed prompts. Later, optimization-based methods Zou et al. (2023); Liu et al. (2024b); Guo et al. (2024) emerged, using iterative optimization strategies to refine harmful prompts and enhance their stealthiness. Model-based methods Chao et al. (2023); Ding et al. (2024); Paulus et al. (2024); Huang et al. (2024) use attacker LLMs to autonomously generate and improve jailbreak prompts. Meanwhile, rule-based methods Jiang et al. (2024); Liu et al. (2024d) rewrite prompts using predefined rules that effectively deceive the model. In this work, we analyze the mechanisms behind these jailbreak techniques.
@@
Defense strategies for jailbreak attacks typically fall into three categories. One common approach enhances model safety and robustness by reformulating inputs to counter jailbreak prompts through techniques such as backtranslation Wang et al. (2024), paraphrasing Jain et al. (2023), self-reminding Xie et al. (2023), and adversarial prompting Zhou et al. (2024a). Another approach focuses on identifying jailbreak prompts using indicators, such as classifying input sequences based on perplexity and sequence length Alon and Kamfonas (2023) or using linear classifiers to detect jailbreak activations and decide whether to respond Zhao et al. (2024a). A more recent trend directly manipulates the model’s internal representations Xu et al. (2024); Li et al. (2024); Liu et al. (2024a) or editing the model Zhao et al. (2024b). Unlike these methods, our approach directly constrains activations within a safety boundary, avoiding extra tokens or modules.
@@
The growing concern about LLM safety has sparked increasing interest in interpreting LLM features in jailbreak prompts. For example, Ball et al. (2024b) identified a common mechanism whereby jailbreaks reduce the harmfulness perception in most LLMs. Similarly, Li et al. (2024) investigated patterns that trigger the model to recognize safety issues. Zhou et al. (2024b) discovered that the vocabulary mappings of activations significantly changed when processing jailbreak inputs. Research efforts have also focused on proposing corresponding defense methods based on interpretability results. For instance, Zhao et al. (2024a) and Shen et al. (2024) modeled how jailbreak activations transfer between benign and harmful activation spaces as layers deepen. They designed adaptive defenses whose strength varies across different layers. However, the limited data in all previous studies often led to controversy and ambiguity.
@@ 3.1 Controversy in Literature
Disputes on layer importance.
There have been different opinions about which layers are most important, mainly due to varying perspectives. For example, He et al. (2024) argued that the low layers were essential because they treated jailbreak activations as benign ones. Zhou et al. (2024b) believed that the low and the middle layers were essential based on tokens generated from activations. Li et al. (2024) argued that deep layers were critical, as their proposed safety pattern showed greater values in these layers. Similarly, He et al. (2024) focused on activations, as we do as well. However, their experimental setup suffers from a key limitation: a small sample size used in the analysis. As shown in Figure 2(a), under the original settings, benign and harmful activations are linearly separable, with most jailbreak samples misclassified as benign. After scaling, half of the jailbreak activations align with harmful ones, making the separation nonlinear. This underscores the risks of small-scale studies and the need for large-scale analysis.
@@
There are differing perspectives on the internal mechanism of jailbreaks. Zhao et al. (2024a) viewed jailbreak as a gradual process, where activations transition from harmful to benign spaces as they pass from the low to the deep layers. In contrast, Shen et al. (2024) argued that jailbreaks manifest abruptly in the deeper layers, with activations positioned between harmful and benign spaces. Both studies relied on fewer than 100 samples per activation type, contributing to similar inconsistencies. As shown in Figure 2(b), before scaling, jailbreak activations align closely with benign ones, consistent with Zhao et al. (2024a). However, after scaling, they diverge from both harmful and benign activations. When considering Shen et al. (2024), similar conflicts arise between pre- and post-scaling results (details in Appendix C.1, Figure 5). These conflicting observations cast doubt on their conclusions regarding the true mechanism of jailbreaks.
Figure 2: PCA visualizations reveal the limitations of small-sample sizes: (a) Top: 100 samples per type He et al. (2024); Bottom: 5,000 benign and harmful samples. (b) Top: 60 samples per type Zhao et al. (2024a); Bottom: 500 samples per type.
@@ 3.2 Experimental Settings
Dataset and Model.
Our dataset consists of 32,507 samples in three categories: benign, harmful, and jailbreak. For benign samples, we used a subset from the Alpaca dataset Taori et al. (2023), as it has been carefully curated to ensure that only safe content is included. We collected harmful samples from five datasets in RedEval Bhardwaj and Poria (2023) and AdvBench Zou et al. (2023). These datasets encompass a wide variety of harmful and misleading prompts. For jailbreak samples, we applied seven different jailbreak attack methods, including GCG-Individual and GCG-Universal Zou et al. (2023), AdvPrompter Paulus et al. (2024), COLD-Suffix Attack Guo et al. (2024), AutoDAN Liu et al. (2024b), PAIR Chao et al. (2023), and ArtPrompt Jiang et al. (2024), on AdvBench. Detailed statistics can be found in Table 4 in Appendix A. We used Vicuna-7B-v1.3 Chiang et al. (2023), a 32-layer Transformer, fine-tuned on a safety-aligned LLM. Although it can identify harmful information, it remains vulnerable to jailbreak attacks Chu et al. (2024), making it an ideal model for analyzing attacks and defenses.
@@
Our projection was based on the activation, the last token vector of the input. As discussed in §3.1, linear classification may be insufficient to model the boundary between different prompts. Therefore, we adopted Multi-Dimensional Scaling (MDS) Carroll and Arabie (1998) as our dimensionality reduction method, as PCA assumes linear separability Anowar et al. (2021), and t-SNE focuses on preserving local structures while distorting the global distribution Van der Maaten and Hinton (2008).
@@ 3.3.1 Activation Distribution
We present the activation distributions of representative layers in Figure 1, with the full version available in Figures LABEL:fig:all_subfig1,LABEL:fig:all_subfig2,LABEL:fig:all_subfig3 in Appendix B. Our direct observations are as follows: (1) Harmful and benign activations overlap and are not linearly separable in most layers. There are no clear linear decision boundaries to distinguish between harmful and benign activations, highlighting the ineffectiveness of treating jailbreak activations as a simple linear classification problem. (2) Jailbreak activations shift significantly out of the harmful activation space, forming distinct regions with minimal overlap. This shift is consistent across different jailbreaks, starting in the low layers and persisting throughout the model, rather than originating in the middle or deep layers Zhou et al. (2024b); Shen et al. (2024). Additionally, the shift is most pronounced in the low and middle layers.
@@ 3.3.2 Safety Boundary
Based on observations of jailbroken and harmful activations, we propose that jailbreak mechanisms are better explained by activations shifting outside a closed safety boundary, rather than simply crossing from one side of a linear boundary to the other. To validate this, we conduct experiments and introduce two key concepts below. Randomized Activation Shifting (RAS): It is an approximation of jailbreak attacks by randomly shifting the original activation. Assume the activation on the $l$-th layer is $a_{l}$, we shift it by a distance $r$ in a random direction $\hat{u}$:
@@
[…]
Most Vulnerable Distance (MVD): It is a measurable approximation of the safety boundary. MVD represents the specific distance at which the DSR experiences the sharpest drop, indicating the point where the model’s ability to detect harmful activations becomes most vulnerable:
@@
where $\frac{\mathrm{d(DSR)}}{\mathrm{d}r}$ denotes the rate of change of DSR with respect to $r$. As shown in Figure 3(b), MVD increases with model depth, indicating that the safety boundary expands in deeper layers and requires larger shifts to bypass safety checks. This aligns with Figure 1: in lower layers (Figure 1(a)), jailbreak activations scatter widely but surpass the safety boundary; in middle layers (Figure 1(b)), jailbreak activations align more with benign activations while remaining outside the boundary; in deeper layers (Figure 1(c)), jailbreak activations cluster, partially overlapping the safety boundary.